OpenStudy (anonymous):

simplify the following radicals please show work 1. √8 * 0.25√6 2. (-√5 +2)(4√3 - √7) 3. 3/√2 - 3

OpenStudy (anonymous):

1. \[\sqrt{8}*0.25*\sqrt{6} = \sqrt{48}/4 = \sqrt{16(3)}/4 = 4\sqrt{3}/4 = \sqrt{3}\] 2. \[(-\sqrt{5}+2)(4\sqrt{3}-\sqrt{7}) = -4\sqrt{15} + \sqrt{35} +8\sqrt{3} - \sqrt{7}\] 3. already 'simple'.